At age 60, there is an equal number of men and women in the population. Schizophrenia is independent of gender* and affects 2% of the population. What is the probability that a randomly picked person from the population is a 60 year old man with schizophrenia?
Answer #1

Given that Schizophrenia is independent of gender. Hence,

P(Randomly picked person from the population is a 60 year old man with schizophrenia)

= P(Man)*P(Schizophrenia)

= 0.50 * 0.02

= 0.01
Answer #2

Given that Schizophrenia is independent of gender. Hence,

P(Randomly picked person from the population is a 60 year old man with schizophrenia)

= P(Man)*P(Schizophrenia)

= 0.50 * 0.02

= 0.01